The shortest leg of a triangle is 7 inches shorter than the other leg. The hypotenuse of this triangle is 13 inches. What are the lengths of the two legs of this triangle?
Answer by ikleyn(53763)

.

The second famous Pythagorean triple after (3,4,5)  with  3^2 + 4^2 = 5^2


is (5,12,13)  with  5^2 + 12^2 = 25 + 144 = 169 = 13^2.


So, your triangle is (5,12,13) inches.    ANSWER



Alternatively,  you can solve this quadratic equation

    x^2 + (x-7)^2 = 13.


You will get the same triangle, then.


After doing some work, of course.

Solved.